Calculate the pressure exerted on an elevator floor which has an area of 6m²

1
Expert's answer
2021-12-30T11:53:38-0500

By the definition of the pressure, we have:


"P=\dfrac{F}{A}=\dfrac{1500\ N}{6\ m^2}=250\ Pa."